# Optimal Control of Behavioral-Feedback SIR Epidemic Model

### v1

@@ 3.3 Dynamic behavior of the uncontrolled BF-SIR epidemic model
Let $\mathcal{I}=\{(x,\bar{y})\in\mathcal{S}\,:\,x\leq\bar{x}\}$. Since $\mathcal{S}$ is positively invariant by Proposition 1 a solution $(x(t),y(t))$ can exit $\mathcal{D}_{\bar{y}}^{-}$ by crossing either $\Gamma^{-}(\bar{x},\bar{y})$ or $\mathcal{I}$. The first option is not possible for the uniqueness of the solution of the ODE. Since $R(\bar{x},\bar{y})=1$ by construction, it follows from (5) that $R(x,y)<1$ for every point $(x,y)\in\mathcal{I}$. Hence, the trajectory cannot leave $\mathcal{D}_{\bar{y}}$ by crossing $\mathcal{I}$. This concludes the proof.
@@
    for every $(x,y)$ in $\mathcal{D}_{\bar{y}}^{+}$, there exists a $\mathcal{C}^{2}$ map $h:\mathcal{D}_{\bar{y}}^{+}\to(\bar{x},1-\bar{y}]$ such that $(h(x,y),\bar{y})$ belongs to $\Gamma(x,y)$ and $h(x,y)\leq x$, where equality holds true if and only if $y=\bar{y}$;
@@
  4. (iv)
Proof 3.19.
(i) Recall from Remark 3.14 that $R(x,y)>1$ for every $(x,y)$ in $\mathcal{D}_{\bar{y}}^{+}$. This implies that every trajectory in $\mathcal{D}_{\bar{y}}^{+}$ is such that $y(t)$ is strictly increasing in $t$ as long as the solution remains in $\mathcal{D}_{\bar{y}}^{+}$. It thus follows from Proposition 3.7(ii) that the solution must necessarily exit $\mathcal{D}_{\bar{y}}^{+}$ in finite time. Since $\mathcal{S}$ is positively invariant by Proposition 1, the trajectory can leave $\mathcal{D}_{\bar{y}}^{+}$ by crossing either $\Gamma^{-}(\bar{x},\bar{y})$ or the set
@@
to be the solution of (14) at time $t$ with initial condition $(x_{0},y_{0})$. Now, note that, since the right hand side of (14) is of class $\mathcal{C}^{2}$, $\phi$ is of class $\mathcal{C}^{2}$ both in $t$ and in $(x_{0},y_{0})$. We now define $T(x_{0},y_{0})$ to be the minimum time $t$ for which $\phi(t,x_{0},y_{0})\in\mathcal{J}$, well defined because of item (i), and define $h:\mathcal{D}_{\bar{y}}^{+}\to(\bar{x},1-\bar{y}]$ by putting
@@
Note that, for every $(x_{0},y_{0})$ in $\mathcal{D}_{\bar{y}}^{+}$, $y(T(x_{0},y_{0}),x_{0},y_{0})=\bar{y}$ by construction and $\dot{y}(T(x_{0},y_{0}),x_{0},y_{0})>0$ because of Remark 3.14. A direct application of the Implicit Function Theorem [rudin1976principles, Theorem 9.18] (on all points of $\mathcal{D}_{\bar{y}}^{+}$ including the boundary segment $\mathcal{J}$) together with the regularity of $\phi$ yields that $T$ is of class $\mathcal{C}^{2}$. Finally, the $\mathcal{C}^{2}$ regularity of $h$ follows from the analogous regularity of $T$ and $\phi$. Note also that $\dot{x}(T(x_{0},y_{0}),x_{0},y_{0})\leq 0$ for every $(x_{0},y_{0})$ in $\mathcal{D}_{\bar{y}}^{+}$ from Proposition 1(i), and together with $x(T(x_{0},y_{0}),x_{0},y_{0})\leq x_{0}$ leads to $h(x_{0},y_{0})\leq x_{0}$.
(iv) The fact that $h$ is a constant of motion of (14) is a consequence of its definition, since the trajectories that start from points that belong to the same negative semi-orbit will hit segment $\mathcal{J}$ in the same point because of the uniqueness of the solutions.
(iii) If $(x_{0},y_{0})=(\bar{x},\bar{y})$, the thesis is straightforward: fixed an arbitrary $\epsilon>0$, for every $(x,y)\in B_{\epsilon}(x_{0},y_{0})\cap\mathcal{D}_{\bar{y}}^{+}$ we have that, thanks to the definition of $h$ and item (ii),
@@
Such a time $T_{\epsilon}$ exists because of the continuity of the solution and it holds that $\epsilon_{y}=\bar{y}-y(T_{\epsilon},x_{0},y_{0})\in(0,\epsilon/2)$. By regularity with respect to the initial condition, there exists $\delta>0$ such that, for every $(x_{\delta},y_{\delta})\in B_{\delta}(x_{0},y_{0})\cap\mathcal{D}_{\bar{y}}^{+}$,
$\phi(t,x_{0},y_{0})-\phi(t,x_{\delta},y_{\delta})||<\epsilon_{y}\quad\forall t\in[0,T_{\epsilon}]$ (24)
This implies that $\phi(t,x_{\delta},y_{\delta})$ belongs to $\mathcal{D}_{\bar{y}}^{+}$ for all $t\in[0,T_{\epsilon}]$. As a consequence of item (ii), we thus have that
@@
Proof 3.22.
(i) It follows from Lemma 3.18(iv) that $\dot{h}=h_{x}\dot{x}+h_{y}\dot{y}=0$ for every $(x,y)$ in $\mathcal{D}_{\bar{y}}^{+}$ under the uncontrolled BF-SIR model (14). Using (14), we then get item (i).
@@
Lemma 3.18 and Lemma 3.21 guarantee the existence of $h(x,y)$ for every $(x,y)$ in $\mathcal{D}_{\bar{y}}^{+}$ under conditions (5)-(6) and characterize $h(x,y)$. Although its existence and regularity is essential to prove our main result, the explicit form of $h(x,y)$ cannot be found unless in some special cases. For instance, the explicit form of $h(x,y)$ may be found for the classical SIR epidemic model and for the BF-SIR epidemic model with feedback mechanism of Example 2.3, due to the fact that the BF-SIR epidemic model with such feedback mechanism and the classical SIR model admit a non-trivial invariant of motion. We refer to [CDC2025], [Miclo.ea:2022] for more details.
@@ 4.1 Candidate value function
In this section, we compute the cost of the candidate optimal control signal $u^{*}(t)$ defined in (11)–(12).
We start with a simple preliminary result that characterizes the dynamics under the optimal control and the associated cost, when the optimal control is not the trivial control signal $u^{*}\equiv 0$. As pointed out in Remark 3.15, if (16) is not verified, or if (16) holds true and $\tilde{y}\leq\bar{y}$ (cf. Lemma 3.9 for the definition of $\tilde{y}$), then $\mathcal{D}_{\bar{y}}^{-}=\mathcal{D}_{\bar{y}}$, so that the optimal control is $u^{*}\equiv 0$ by Lemma 3.16. Hence, in the next part we shall work under the assumption that (16) holds true and $\tilde{y}>\bar{y}$. We let $T_{0}=T(x_{0},y_{0})$ be the first time the solution of (14) with initial condition $(x_{0},y_{0})$ in $\mathcal{D}^{+}_{\bar{y}}$ hits the threshold $\bar{y}$ and let
@@
Then the corresponding solution of the CBF-SIR epidemic model (4) $(x^{*}(t),y^{*}(t))$ has the following properties:
@@
We observe that the vector function in (33) coincides with $(x^{*}(T_{0}),y^{*}(T_{0}))$ for $t=T_{0}$. Moreover a direct check shows that such a vector function is a solution of (4) with control $u^{*}$ in $T_{0},T_{1}]$. This proves item (i). Item (ii) is a direct consequence of the definition of $\mu$ and of item (i). To prove item (iii) notice that, by construction the inequality $y^{*}(t)\leq\bar{y}$ is satisfied till time $T_{1}$. Notice now that the solution $(x^{*}(t),y^{*}(t))$ for $t>T_{1}$ coincides with the solution of the uncontrolled model with initial condition $x^{*}(T_{1})=\bar{x}$ and $y^{*}(T_{1})=\bar{y}$. Since $R(\bar{x},\bar{y})=1$, it follows from Proposition 3.7(i) that $y^{*}(t)$ is strictly decreasing from this time on. This yields $y^{*}(t){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}<}\bar{y}$ also for $t>T_{1}$ and proves item (iii). Finally,
@@
Proof 4.27.
(i) Note that $V(x,y)=0$ for every $(x,y)$ in $\Gamma^{-}(\bar{x},\bar{y})$ since $\Gamma^{-}(\bar{x},\bar{y})$ belongs to $\mathcal{D}_{\bar{y}}^{-}$. Moreover, it follows from Lemma 3.18(iii) and from the Definition of $V$ that for every $(x,y)$ in $\Gamma^{-}(\bar{x},\bar{y})$
@@
This, together with the continuity of $h(x,y)$ established in Lemma 3.18(i), proves continuity of $V$ on $\mathcal{D}_{\bar{y}}$. The differentiability of $V$ on $\mathcal{D}_{\bar{y}}^{+}$ follows from the differentiability of $h$ established in Lemma 3.18(ii).
@@
Lemma 4.26 states that the candidate value function $V(x,y)$ is continuous on $\mathcal{D}_{\bar{y}}$, but continuously differentiable on $\mathcal{D}_{\bar{y}}^{+}$ only. In fact, it is not hard to prove that $V(x,y)$ is not continuously differentiable on $\Gamma^{-}(\bar{x},\bar{y})$. For this reason, we cannot apply the Hamilton-Jacobi-Bellman equation [liberzon, SCC.Kamalapurkar.Walters.ea2018] to prove that $u^{*}$ is the optimal control. In the next section we shall provide an alternative prove for our main result.
@@
If $(x_{0},y_{0})$ belongs to $\mathcal{D}_{\bar{y}}^{-}$, the uncontrolled solution remains below $\bar{y}$ at all times thanks to Lemma 3.16. Therefore $u\equiv 0$ is an optimal control that satisfies Theorem 2.5.
@@
where the second equality follows from (36). Because of Definition (10) and Assumption 1, the function $\rho(x,y)$ is increasing in $x$ and decreasing in $y$. Since $h(x^{\circ}(t),x^{\circ}(t))\leq x^{\circ}(t)$ by Lemma 3.18(ii) and $y^{\circ}(t)\leq\bar{y}$ for every $t\leq t^{\circ}$, we thus have
@@
where the first inequality follows from (41)-(42) and from the left-hand inequality of Lemma 4.26(iii), the equality from (4), the second inequality from the right-hand inequality of Lemma 4.26(iii) and the last one from the feasibility of $u^{\circ}$. On the other hand, (34) implies that
@@
where the first inequality follows from the monotonicity of $x^{\circ}(t)$ and the non-negativity of $V_{x}$, the equality from (4), the second inequality from the left-hand inequality of Lemma 4.26(iii) and from $(x^{\circ},y^{\circ})\in\mathcal{D}_{\bar{y}}^{+}$, which together with Remark 3.15 imply that $R(x^{\circ},y^{\circ})(1-u^{\circ})-1\leq R(x^{\circ},y^{\circ})$, the third one from the right-hand inequality of Lemma 4.26(iii) and the last one from Assumption 1. Inequalities (43) and (45) together with relation (44) yield that $\dot{V}(x^{\circ}(t),y^{\circ}(t))$ is a bounded function on the complementary of a negligible set. Consequently, $V(x^{\circ}(t),y^{\circ}(t))$ is an absolutely continuous function on $0,+\infty)$ and the fundamental theorem of calculus holds. From relations (43) and (44) we thus obtain
@@
Consider now the solution of the candidate optimal control $u^{*}$ and the associated solution $(x^{*}(t),y^{*}(t))$ with initial condition $(x_{0},y_{0})$. From Proposition 4.24(i), we know that the solution belongs to $\mathcal{D}_{\bar{y}}^{+}$ for $t$ in $T_{0},T_{1})$. For $t$ in $T_{0},T_{1})$, we have the following chain of equalities:
@@
where the last equality follows from Remark 4.28 and the last inequality from (46). Finally, relations (40) and (48) imply
@@
As $u$ was chosen to be any feasible control function, this proves the optimality of $u^{*}$. Since $u^{*}$ is feasible by Proposition 4.24(iii), the proof of existence in Theorem 2.5 is now completed.
We now prove uniqueness. Suppose $u$ in $\mathcal{U}_{x_{0},y_{0}}^{\bar{y}}$ is optimal. Following previous considerations, similarly defining $t^{o}$ as the first time the corresponding solution $(x^{u},y^{u})$ exits the region $\mathcal{D}_{\bar{y}}^{+}$ and defining $u^{o}$ as in (39), we notice that inequality (40) must now be an equality. This forces $u(t)=0$ for $t>t^{o}$. We now consider $u(t)$ for $t\leq t^{o}$ and we prove that $u$ must be 0 before the solution reaches the threshold value $\bar{y}$. We define $\mathcal{C}=\{t\in[0,t^{o}]\,|\,y^{u}(t)<\bar{y},u(t)>0\}$ and prove that, indicating with $\mu$ the Lebesgue measure, $\mu(\mathcal{C})=0$. Suppose by contradiction that instead $\mu(\mathcal{C})>0$. Then standard considerations on the Lebesgue integral imply that
@@
$\int_{0}^{t^{\circ}}u(s)\mathrm{d}s>\int_{0}^{t^{o}}\frac{y^{u}(s)}{\bar{y}}u^{\circ}(s)\mathrm{d}s\geq V(x_{0},y_{0})\,.$ (49)
This contradicts the supposed optimality of $u$ and implies that $\mu(\mathcal{C})=0$. Since $u$ is assumed to be right-continuous and $y$ is continuous, this implies that $u(t)=0$ for every $t$ such that $y^{u}(t)<\bar{y}$. In conclusion, $u$ has to be null as long as the second component $y^{u}(t)$ of the solution stays below the threshold value $\bar{y}$ and from the first time the solution exits $\mathcal{D}_{\bar{y}}^{+}$. Therefore, it can only be non null when the solution is such that $y^{u}(t)=\bar{y}$ and $x^{u}(t)>\bar{x}$. We now conclude by showing that, necessarily $u(t)=\rho(x^{u}(t),\bar{y})$ in all such points. We first notice that we must have $u(t)\geq\rho(x^{u}(t),\bar{y})$. Indeed, if $u(\bar{t})<\rho(x^{u}(\bar{t}),\bar{y})$ for some $\bar{t}$, equations 4 and the right-continuity of $u$ imply that $u$ would not be feasible as it would allow the component $y^{u}$ to exceed $\bar{y}$ in a right neighborhood of $\bar{t}$. Moreover, if $u(\bar{t})>\rho(x^{u}(\bar{t}),\bar{y})$ at some point $\bar{t}$, then, similarly, $y^{u}$ would be decreasing in a right neighborhood of $\bar{t}$ and thus $u$ would be non null in instants where $y^{u}$ is below $\bar{y}$. As this fact was already shown to lead to non optimality, is also forbidden for $u$. This proves that $u(t)=\rho(x^{u}(t),\bar{y})$ the solution is such that $y^{u}(t)=\bar{y}$ and $x^{u}(t)>\bar{x}$, and thus is necessarily identical to $u^{*}$.

### v2

@@ 3.3 Dynamic behavior of the uncontrolled BF-SIR epidemic model
Let $\mathcal{I}=\{(x,\bar{y})\in\mathcal{S}\,:\,x\leq\bar{x}\}$. Since $\mathcal{S}$ is positively invariant by Proposition 1 a solution $(x(t),y(t))$ can exit $\mathcal{D}_{\bar{y}}^{-}$ by crossing either $\Gamma^{-}(\bar{x},\bar{y})$ or $\mathcal{I}$. The first option is not possible for the uniqueness of the solution of the ODE. Since $R(\bar{x},\bar{y})=1$ by construction, it follows from (5) that $R(x,y)<1$ and thus $\dot{y}\leq 0$ for every point $(x,y)\in\mathcal{I}$. Hence, the trajectory cannot leave $\mathcal{D}_{\bar{y}}$ by crossing $\mathcal{I}$. This concludes the proof.
@@
    for every $(x,y)$ in $\mathcal{D}_{\bar{y}}^{+}$, there exists a $\mathcal{C}^{2}$ map $h:\mathcal{D}_{\bar{y}}^{+}\to(\bar{x},1-\bar{y}]$ such that $(h(x,y),\bar{y})$ belongs to $\Gamma^{+}(x,y)$ and $h(x,y)\leq x$, where equality holds true if and only if $y=\bar{y}$;
@@
Remark 3.19.
Proof (of Lemma 3.18): (i) Recall from Remark 3.14 that $R(x,y)>1$ for every $(x,y)$ in $\mathcal{D}_{\bar{y}}^{+}$. This implies that every trajectory in $\mathcal{D}_{\bar{y}}^{+}$ is such that $y(t)$ is strictly increasing in $t$ as long as the solution remains in $\mathcal{D}_{\bar{y}}^{+}$. It thus follows from Proposition 3.7(ii) that the solution must necessarily exit $\mathcal{D}_{\bar{y}}^{+}$ in finite time. Since $\mathcal{S}$ is positively invariant by Proposition 1, the trajectory can leave $\mathcal{D}_{\bar{y}}^{+}$ by crossing either $\Gamma^{-}(\bar{x},\bar{y})$ or the set
@@
be the solution of (14) at time $t$ with initial condition $(x_{0},y_{0})$. Now, note that, since the right hand side of (14) is of class $\mathcal{C}^{2}$, $\phi$ is of class $\mathcal{C}^{2}$ both in $t$ and in $(x_{0},y_{0})$ [21, Chapter 5]. We now define $T(x_{0},y_{0})$ to be the minimum time $t$ for which $\phi(t,x_{0},y_{0})\in\mathcal{J}$, well defined because of item (i), and define $h:\mathcal{D}_{\bar{y}}^{+}\to(\bar{x},1-\bar{y}]$ by putting
@@
Note that, for every $(x_{0},y_{0})$ in $\mathcal{D}_{\bar{y}}^{+}$, $y(T(x_{0},y_{0}),x_{0},y_{0})=\bar{y}$ by construction and $\dot{y}(T(x_{0},y_{0}),x_{0},y_{0})>0$ because of Remark 3.14. A direct application of the Implicit Function Theorem [40, Theorem 9.18] (on all points of $\mathcal{D}_{\bar{y}}^{+}$ including the boundary segment $\mathcal{J}$) together with the regularity of $\phi$ yields that $T$ is of class $\mathcal{C}^{2}$ in $(x_{0},y_{0})$. Finally, the $\mathcal{C}^{2}$ regularity of $h$ follows from the analogous regularity of $T$ and $\phi$. Note also that by Proposition 1(i), $h(x_{0},y_{0})=x(T(x_{0},y_{0}),x_{0},y_{0})\leq x_{0}$.
(iii) The fact that $h$ is a constant of motion of (14) is a consequence of its definition, since the trajectories that start from points that belong to the same negative semi-orbit will hit segment $\mathcal{J}$ in the same point because of the uniqueness of the solutions.
(iv) If $(x_{0},y_{0})=(\bar{x},\bar{y})$, the thesis is straightforward: fixed an arbitrary $\epsilon>0$, for every $(x,y)\in B_{\epsilon}(x_{0},y_{0})\cap\mathcal{D}_{\bar{y}}^{+}$ we have that, thanks to the definition of $h$ and item (ii),
@@
Such a time $T_{\epsilon}$ exists because of the continuity of the solution. Let $\epsilon_{y}:=\bar{y}-y(T_{\epsilon},x_{0},y_{0})\in(0,\epsilon/2)$. By regularity with respect to the initial condition [21, Chapter 5], there exists $\delta>0$ such that, for every $(x_{\delta},y_{\delta})\in B_{\delta}(x_{0},y_{0})\cap\mathcal{D}_{\bar{y}}^{+}$,
$\phi(t,x_{0},y_{0})-\phi(t,x_{\delta},y_{\delta})||<\epsilon_{y}\quad\forall t\in[0,T_{\epsilon}].$ (24)
This implies that $\phi(t,x_{\delta},y_{\delta})$ belongs to $\mathcal{D}_{\bar{y}}^{+}$ for all $t\in[0,T_{\epsilon}]$ since $\bar{y}-y(T_{\epsilon},x_{0},y_{0})=\epsilon_{y}$ and $|y(T_{\epsilon},x_{0},y_{0})-y(T_{\epsilon},x_{\delta},y_{\delta})|<\epsilon_{y}$ imply $y(T_{\epsilon},x_{\delta},y_{\delta})<\bar{y}$. As a consequence of item (ii), we thus have that
@@
Proof 3.21.
(i) It follows from Lemma 3.18(iii) that $\dot{h}=h_{x}\dot{x}+h_{y}\dot{y}=0$ for every $(x,y)$ in $\mathcal{D}_{\bar{y}}^{+}$ under the uncontrolled BF-SIR model (14). Using (14), we then get item (i).
@@
Lemma 3.18 and Lemma 3.20 guarantee the existence of $h(x,y)$ for every $(x,y)$ in $\mathcal{D}_{\bar{y}}^{+}$ under conditions (5)-(6) and characterize $h(x,y)$. Although its existence and regularity is essential to prove our main result, the explicit form of $h(x,y)$ cannot be found unless in some special cases. For instance, the explicit form of $h(x,y)$ may be found for the classical SIR epidemic model and for the BF-SIR epidemic model with feedback mechanism of Example 2.3, due to the fact that the BF-SIR epidemic model with such feedback mechanism and the classical SIR model admit a non-trivial invariant of motion. We refer to [3], [35] for more details.
@@ 4.1 Candidate value function
We start by computing the cost of the candidate optimal control signal $u^{*}(t)$ defined in (11)–(12).
As pointed out in Remark 3.15, if (16) is not verified, or if (16) holds true and $\tilde{y}\leq\bar{y}$ (cf. Lemma 3.9 for the definition of $\tilde{y}$), then $\mathcal{D}_{\bar{y}}^{-}=\mathcal{D}_{\bar{y}}$, so that the optimal control is $u^{*}\equiv 0$ by Lemma 3.16. Hence, in the next part we shall work under the assumption that (16) holds true and $\tilde{y}>\bar{y}$. We let $T_{0}=T(x_{0},y_{0})$ be the first time the solution of (14) with initial condition $(x_{0},y_{0})$ in $\mathcal{D}^{+}_{\bar{y}}$ hits the threshold $\bar{y}$ and let
@@
The corresponding solution of the CBF-SIR epidemic model (4) $(x^{*}(t),y^{*}(t))$ has the following properties:
@@
We observe that the vector function in (33) coincides with $(x^{*}(T_{0}),y^{*}(T_{0}))$ for $t=T_{0}$. Moreover a direct check shows that such a vector function is a solution of (4) with control $u^{*}$ in $T_{0},T_{1}]$. This proves item (i). Item (ii) is a direct consequence of the definition of $\mu$ and of item (i). To prove item (iii) notice that, by construction, the inequality $y^{*}(t)\leq\bar{y}$ is satisfied till time $T_{1}$ and the solution $(x^{*}(t),y^{*}(t))$ for $t>T_{1}$ coincides with the solution of the uncontrolled model with initial condition $x^{*}(T_{1})=\bar{x}$ and $y^{*}(T_{1})=\bar{y}$. Since $R(\bar{x},\bar{y})=1$, it follows from Proposition 3.7(i) that $y^{*}(t)$ is strictly decreasing from this time on. This yields $y^{*}(t){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}<}\bar{y}$ for $t>T_{1}$ and proves item (iii). Finally,
@@
Proof 4.26.
(i) Note that $V(x,y)=0$ for every $(x,y)$ in $\Gamma^{-}(\bar{x},\bar{y})$ since $\Gamma^{-}(\bar{x},\bar{y})$ belongs to $\mathcal{D}_{\bar{y}}^{-}$. Moreover, it follows from Lemma 3.18(iv) and from the definition of $V$ that for every $(x,y)$ in $\Gamma^{-}(\bar{x},\bar{y})$
@@
This, together with the continuity of $h(x,y)$ established in Lemma 3.18(ii), proves continuity of $V$ on $\mathcal{D}_{\bar{y}}$. The differentiability of $V$ on $\mathcal{D}_{\bar{y}}^{+}$ follows from the differentiability of $h$ established in Lemma 3.18(ii).
@@
Lemma 4.25 states that the candidate value function $V(x,y)$ is continuous on $\mathcal{D}_{\bar{y}}$, but continuously differentiable on $\mathcal{D}_{\bar{y}}^{+}$ only. In fact, it is not hard to prove that $V(x,y)$ is not continuously differentiable on $\Gamma^{-}(\bar{x},\bar{y})$. For this reason, we cannot apply the Hamilton-Jacobi-Bellman equation [31, 25] to prove that $u^{*}$ is the optimal control. In the next section we shall provide an alternative prove for our main result.
@@
If $(x_{0},y_{0})$ belongs to $\mathcal{D}_{\bar{y}}^{-}$, the uncontrolled solution remains below $\bar{y}$ at all times thanks to Lemma 3.16. Therefore $u\equiv 0$ is the unique optimal control that satisfies Theorem 2.5.
@@
where the second equality follows from (36). Because of Definition (10) and Assumption 1, the function $\rho(x,y)$ is increasing in $x$ and non-increasing in $y$. Since $h(x^{\circ}(t),y^{\circ}(t))\leq x^{\circ}(t)$ by Lemma 3.18(ii) and $y^{\circ}(t)\leq\bar{y}$ for every $t\leq t^{\circ}$, we thus have
@@
where the first inequality follows from (41)-(42), $\dot{x}\leq 0$ and from the left-hand inequality of Lemma 4.25(iii), the equality from (4), the second inequality from the right-hand inequality of Lemma 4.25(iii) and the last one from the feasibility of $u^{\circ}$. On the other hand, (34) implies that
@@
where the first inequality follows from the monotonicity of $x^{\circ}(t)$ and the non-negativity of $V_{x}$, the equality from (4), the second inequality from the left-hand inequality of Lemma 4.25(iii) and from $R(x^{\circ},y^{\circ})(1-u^{\circ})-1\leq R(x^{\circ},y^{\circ})$, the third one from the right-hand inequality of Lemma 4.25(iii) and the last one from Assumption 1. Inequalities (43) and (45) together with relation (44) yield that $\dot{V}(x^{\circ}(t),y^{\circ}(t))$ is a bounded function on the complementary of a negligible set. Consequently, $V(x^{\circ}(t),y^{\circ}(t))$ is an absolutely continuous function on $0,+\infty)$ and the fundamental theorem of calculus holds. From relations (43) and (44) we thus obtain
@@
Consider now the solution of the candidate optimal control $u^{*}$ and the associated solution $(x^{*}(t),y^{*}(t))$ with initial condition $(x_{0},y_{0})$. From Proposition 4.23(i), we know that the solution belongs to $\mathcal{J}\subseteq\mathcal{D}_{\bar{y}}^{+}$ for $t$ in $T_{0},T_{1})$. For $t$ in $T_{0},T_{1})$, we have the following chain of equalities:
@@
where the last equality follows from Remark 4.27 and the last inequality from (46) since $u^{*}=0$ for all $t\in[0,T_{0}]$. Finally, relations (40) and (48) imply
@@
As $u$ was chosen to be any feasible control function, this proves the optimality of $u^{*}$. Since $u^{*}$ is feasible by Proposition 4.23(iii), the proof of existence in Theorem 2.5 is completed.
We next prove uniqueness. Suppose $u$ in $\mathcal{U}_{x_{0},y_{0}}^{\bar{y}}$ is optimal. Following previous considerations, similarly defining $t^{o}$ as the first time the corresponding solution $(x^{u},y^{u})$ exits the region $\mathcal{D}_{\bar{y}}^{+}$ and defining $u^{o}$ as in (39), we notice that inequality (40) must now be an equality. This forces $u(t)=0$ for $t>t^{o}$. We now consider $u(t)$ for $t\leq t^{o}$ and we prove that $u$ must be 0 before the solution reaches the threshold value $\bar{y}$. We define $\mathcal{C}=\{t\in[0,t^{o}]\,|\,y^{u}(t)<\bar{y},u(t)>0\}$ and prove that, indicating with $\mu$ the Lebesgue measure, $\mu(\mathcal{C})=0$. Suppose by contradiction that $\mu(\mathcal{C})>0$. Then standard considerations on the Lebesgue integral imply that
@@
$J(u)\!=\!\int_{0}^{t^{\circ}}u(s)\mathrm{d}s\!>\!\int_{0}^{t^{o}}\frac{y^{u}(s)}{\bar{y}}u^{\circ}(s)\mathrm{d}s\geq V(x_{0},y_{0})\!=\!J(u^{*}).$ (49)
This contradicts the supposed optimality of $u$ and implies that $\mu(\mathcal{C})=0$. Since $u$ is assumed to be right-continuous and $y$ is continuous, this implies that $u(t)=0$ for every $t$ such that $y^{u}(t)<\bar{y}$. In conclusion, $u$ has to be null as long as the second component $y^{u}(t)$ of the solution stays below the threshold value $\bar{y}$ and from the first time the solution exits $\mathcal{D}_{\bar{y}}^{+}$. Therefore, it can only be non null when the solution is such that $y^{u}(t)=\bar{y}$ and $x^{u}(t)>\bar{x}$. We now conclude by showing that, necessarily $u(t)=\rho(x^{u}(t),\bar{y})$ in all such points. We first notice that we must have $u(t)\geq\rho(x^{u}(t),\bar{y})$. Indeed, if $u(\bar{t})<\rho(x^{u}(\bar{t}),\bar{y})$ for some $\bar{t}$, equations (4) and the right-continuity of $u$ imply that $u$ would not be feasible as it would allow the component $y^{u}$ to exceed $\bar{y}$ in a right neighborhood of $\bar{t}$. Moreover, if $u(\bar{t})>\rho(x^{u}(\bar{t}),\bar{y})$ at some point $\bar{t}$, then, similarly, $y^{u}$ would be decreasing in a right neighborhood of $\bar{t}$ and thus $u$ would be non null in instants where $y^{u}$ is below $\bar{y}$. As this fact was already shown to lead to non optimality, it is also forbidden for $u$. This proves that $u(t)=\rho(x^{u}(t),\bar{y})$ whenever the solution is such that $y^{u}(t)=\bar{y}$ and $x^{u}(t)>\bar{x}$, and thus is necessarily identical to $u^{*}$.